OpenStudy (puertoricangirl):

The number of chips of different colors in Rick's bag is shown below: •3 red chips •11 green chips •7 yellow chips Rick takes out a chip from the bag randomly without looking. He replaces the chip and then takes out another chip from the bag. What is the probability that Rick takes out a yellow chip in both draws? A. 7 over 21 multiplied by 6 over 20 equal to 42 over 420 B. 7 over 21 plus 6 over 20 equal to 266 over 420 C.7 over 21 multiplied by 7 over 21 equal to 49 over 441 D. 7 over 21 plus 7 over 21 equal to 14 over 21

Mehek (mehek14):

\(\dfrac{yellow}{total}=\dfrac{7}{21}\) Since he replaces it, the amount stays the same \(\dfrac{7}{21}*\dfrac{7}{21}=\)

OpenStudy (puertoricangirl):

so it's C

OpenStudy (puertoricangirl):

@Mehek14 @supercalifragalisticexspeaalli